How many possible samples does it have when a population consist of 12 and you have to get a sample size of 4 from it. *

1
Expert's answer
2022-04-12T05:04:14-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! };\\
\begin{pmatrix}
 12 \\
 4
\end{pmatrix}=\cfrac{12! } {4! \cdot8! }=\cfrac{9\cdot10\cdot11\cdot12}{2\cdot3\cdot4}=495."